OpenStudy (anonymous):

use mathematical induction to prove n sigma i=1 1/i(i+1) = n/n+1

OpenStudy (anonymous):

posting a solution in a sec, how familiar are you with proofs by induction?

OpenStudy (anonymous):

slightly familiar i know 3 steps i think im just having trouble moving forward from k sigma i = 1 1/i(i+1) + 1/ (k+1)(k+2)

OpenStudy (anonymous):

alright, so you added the "missing term" \[\frac{1}{(k+1)(k+2)}\] to your summation, and you want to simplify? just trying to make sure we are the same page since there are a couple of ways to tackle this problem.

OpenStudy (anonymous):

yes, i want to simplify

OpenStudy (anonymous):

by the way, this is my solution, but I would still like to do it with your approach.

OpenStudy (anonymous):

alright so you have: \[\frac{1}{(k+1)(k+2)}+\sum_{i=1}^{k}\frac{1}{i(i+1)} \] to simply, use your inductive hypothesis, which is that the summation equals k/(k+1). Then we have: \[\frac{1}{(k+1)(k+2)}+\frac{k}{k+1}\] and the game is to simplify as much as we can. i believe that work is done in my 3rd pic.

OpenStudy (anonymous):

but how do i know the summation is k/k+1 and not 1/k+1?

OpenStudy (anonymous):

because we are trying to prove: \[\sum_{i=1}^{n}\frac{1}{i(i+1)} = \frac{n}{n+1}\] We assume this is true for some fixed natural number 'k'. so we have: \[\sum_{i=1}^{k}\frac{1}{i(i+1)} = \frac{k}{k+1}\] that is why we use k/(k+1), not 1/(k+1)

OpenStudy (anonymous):

gotta use the formula you are trying to prove.

OpenStudy (anonymous):

oh... right.

OpenStudy (anonymous):

you also have to know what you are aiming for as well. hopefully, those fractions should simplify to: \[\sum_{i=1}^{k+1}\frac{1}{i(i+1)}= \frac{(k+1)}{(k+1)+1} = \frac{k+1}{k+2}\]

OpenStudy (anonymous):

ok now im seeing it...

OpenStudy (anonymous):

sweet :) induction is a powerful tool, but getting the hang of it can be rough sometimes (especially with inequalities, they can die in a fire >.<)

OpenStudy (anonymous):

so 1/k+1 + k/k+1 = k+1/K+1?

OpenStudy (anonymous):

that last kkay is not meant to be a capital

OpenStudy (anonymous):

you want to show that: \[\frac{1}{(k+1)(k+2)} + \frac{k}{k+1} \Rightarrow \frac{k+1}{k+2}\] somehow, by combining fractions, simplifying, etc.

OpenStudy (anonymous):

yeah, im just looking at combining the denominator... because on the numerator, i can see that its just 1+k... but on the bottom does (k+1)(k+2) + k+1 just combine to (k+1)(k+2) ?

OpenStudy (anonymous):

you gotta get a common denominator before you add those fractions up: \[\frac{1}{(k+1)(k+2)}+\frac{k}{k+1} = \frac{1}{(k+1)(k+2)} +\frac{k(k+2)}{(k+1)(k+2)}\] \[= \frac{1+k(k+2)}{(k+1)(k+2)}\]

OpenStudy (anonymous):

oh damn

OpenStudy (anonymous):

the rest of the solution is in the 3rd pic (except i used the dummy variable n, instead of k)

OpenStudy (anonymous):

ah eys now on to the next induction :)